OpenStudy (anonymous):

Solve, x-(3/5) = (7/10 x minus 3 over 5 equals 7 over 10

OpenStudy (anonymous):

Multiply both sides by 10, you get: \[10x-6=7 \implies 10x=13 \implies x={13 \over 10}\]

OpenStudy (anonymous):

@AnwarA Where did you get the 6 from?

OpenStudy (anonymous):

\[-\frac{3}{5}\times 10=-6\]

OpenStudy (anonymous):

As satellite said.

OpenStudy (anonymous):

The answer in the book says its negative 1 over 10.

OpenStudy (anonymous):

I wanna know the steps to solve this equation...

OpenStudy (anonymous):

is it \[x-\frac{3}{5}=\frac{7}{10}\]

OpenStudy (anonymous):

or \[\frac{x-3}{5}=\frac{7}{10}\]?

OpenStudy (anonymous):

yes, the first one

OpenStudy (anonymous):

well then the answer in the book is wrong, because anwar is right.

OpenStudy (anonymous):

one step problem if you do not mind the fractions: \[x-\frac{3}{5}=\frac{7}{10}\] add \[\frac{3}{5}\] to both sides to get \[x=\frac{7}{10}+\frac{3}{5}=\frac{7}{10}+\frac{6}{10}=\frac{13}{10}\]

OpenStudy (anonymous):

No i figured it out. 3/5 is equal to 6/10 so to solve for x i have to get x alone so i minus 6/10 from both sides & 7/10 minus 6/10 equals 1/10 so that makes x=1/10

OpenStudy (gw2011):

3/5 does equal 6/10 but you have to add 6/10 to both sides and not subtract it because you have a -3/5 and not a +3/5

OpenStudy (anonymous):

unfortunately \[-\frac{3}{5}-\frac{3}{5}=-\frac{6}{5}\] not 0. you have to add \[\frac{3}{5}\] to both sides

OpenStudy (anonymous):

scratch that, idk why it says 1/10 for the answer in the book. i'll leave that for the teacher to explain.thanks anyways :)

OpenStudy (radar):

x-3 7 ---- =---- multiply through by 10 5 10 2(x-3) = 7 2x-6 = 7 2x=13 x=13/2